# Reasoning or Memorization? Unreliable Results of Reinforcement Learning Due to Data Contamination

### v2

@@ Abstract
Reasoning in large language models has long been a central research focus, and recent studies employing reinforcement learning (RL) have introduced diverse methods that yield substantial performance gains with minimal or even no external supervision. Surprisingly, some studies even suggest that random or incorrect reward signals can enhance performance. However, these breakthroughs are predominantly observed for the mathematically strong Qwen2.5 series on benchmarks such as MATH-500, AMC, and AIME, and seldom transfer to models like Llama, which warrants a more in-depth investigation. In this work, our empirical analysis reveals that pre-training on massive web-scale corpora leaves Qwen2.5 susceptible to data contamination in widely used benchmarks. Consequently, conclusions derived from contaminated benchmarks on Qwen2.5 series may be unreliable. To obtain trustworthy evaluation results, we introduce a generator that creates fully clean arithmetic problems of arbitrary length and difficulty, dubbed RandomCalculation. Using this leakage-free dataset, we show that only accurate reward signals yield steady improvements that surpass the base model’s performance boundary in mathematical reasoning, whereas random or incorrect rewards do not. Moreover, we conduct more fine-grained analyses to elucidate the factors underlying the different performance observed on the MATH-500 and RandomCalculation benchmarks. Consequently, we recommend that future studies evaluate models on uncontaminated benchmarks and, when feasible, test various model series to ensure trustworthy conclusions about RL and related methods. Code is available at https://github.com/wumingqi/LLM-Math-Evaluation.
@@ 1 Introduction
In recent years, advances in reinforcement learning (RL) have markedly strengthened the reasoning abilities of large language models (LLMs). Flagship systems, including OpenAI’s o1 (OpenAI, 2024b; 2025; a), DeepSeek-R1 (DeepSeek-AI et al., 2025), and QwQ (Qwen Team, 2024; 2025), already match or exceed human-level accuracy on a variety of challenging benchmarks. Among open-source contenders that excel in these mathematical benchmarks, the Qwen family (Yang et al., 2024a; b; 2025), ranging from 0.5B to 72B and pre-trained on up to 36T high-quality tokens, produces state-of-the-art results in language understanding, mathematics, programming, and preference alignment.
Within this landscape, mathematical reasoning emerges as a particularly discriminative test bed because it demands precise symbolic manipulation and multi-step logical deduction. Standard suites such as MATH-500 (Hendrycks et al., 2021), AIME (Li et al., 2024), AMC (Li et al., 2024), and Minerva Math (Lewkowycz et al., 2022) require models to parse natural-language problem statements, uncover the latent mathematical structure, and generate exact numeric answers. Recent work has further enhanced this capability by reinforcement learning with verifiable rewards (RLVR) (DeepSeek-AI et al., 2025): rule-based reward that returns 1 when the predicted answer equals ground truth and 0 otherwise. Because the reward is computed analytically, RLVR removes the need for a separate learned reward model, lowering computational cost while providing a precise training signal, especially attractive for domains like mathematics, where solutions are unambiguous.
@@
Qwen2.5-7B Output it true that $2^{24}$ is a perfect $n^{\text{th}}$ power?
To determine for how many positive integers $n>1$ the number $2^{24}$ is a perfect $n$-th power… First, we need to find… Let’s count these values. There are 77 such values… The final answer is $\boxed{7}$.
Llama3.1-8B Output the set of positive divisors of $n$ (including 11 and $n$) less than 5050 ? (A) 15 (B) 16 (C) 17 (D) 18 (E) 19
The number of positive divisors of $n$ is given by… The number of factors of each of these numbers is… The answer is $\boxed{\textbf{(B) }16}$
@@
Although RL nominally depends on accurate reward signals to guide training, recent studies (Zuo et al., 2025; Shao et al., 2025) find that even random or incorrect rewards can improve Qwen’s performance on standard math benchmarks, while the same procedures offer little or no benefit to Llama3.1-8B (Dubey et al., 2024). To understand why these seemingly problem-agnostic rewards help Qwen but not Llama, we undertake a systematic comparison of two model families under identical training protocols. We consider two working hypotheses to explain this phenomenon. (i) Data Contamination: Considering Qwen2.5 is pre-trained on massive web-scale corpora from the Internet, including GitHub repositories that store benchmark problems alongside their official solutions. If segments of the evaluation benchmarks leaked into the pre-training corpus, spurious rewards could cue the model to retrieve memorized answers rather than acquire new reasoning skills. (ii) Strong Math Capacity: Qwen’s pre-training endows it with better mathematical capacity than Llama, so even noisy policy-gradient updates appear to help on MATH-500. However, if strong capacity is the real driver, the same spurious rewards should still work on a clean benchmark. Distinguishing between these possibilities requires both a leakage audit and a rigorously out-of-distribution RLVR evaluation.
To assess the extent of potential data contamination in popular mathematical benchmarks, we propose two metrics: partial-prompt completion rate (can the model reconstruct the tail of a problem?) and partial-prompt answer accuracy (can the model give the correct answer with an incomplete problem?). As shown in Fig. 1, Qwen can indeed complete the problem accurately and provide the correct answer, whereas Llama fails. Furthermore, prompting with the first 60% of each MATH-500 problem, Qwen2.5-Math-7B regenerates the remaining 40% with a 54.60% exact-match rate and answers 53.6% of these incomplete problems correctly. In contrast, Llama3.1-8B scores 3.8% and 2.4% on both metrics. Crucially, on the newly released LiveMathBench (version 202505) (Liu et al., 2024), Qwen’s completion rate drops sharply to 0.0%, consistent with Llama’s 0.0%. Its partial-prompt answer accuracy also falls to just 2.0%, comparable to Llama’s 1.0%. These results confirm that Qwen’s pre-training corpus suffers from test data contamination. So, results derived from MATH-500 and similar datasets for Qwen should be interpreted with caution.
@@ 2 Related Works
A growing body of work investigates how reinforcement learning (RL) can amplify the mathematical-reasoning capacity of the open-source Qwen2.5 family. Early studies use verifiable rewards that score an answer as 1 / 0 by exact numerical agreement. Test-time RL (Zuo et al., 2025) applies this signal on-the-fly during inference and yields sizeable gains on MATH-500 and AIME2024. Subsequent efforts pursue extreme data efficiency: few labeled (Wang et al., 2025a; Li et al., 2025) or even unlabeled examples (Gao et al., 2025; Zhao et al., 2025a) can suffice to boost performance. A parallel research replaces external supervision with intrinsic signals derived from the model itself: Prabhudesai et al. (2025); Agarwal et al. (2025) reward low-entropy output distributions, while Zhao et al. (2025b); Shafayat et al. (2025) rely on self-consistency or self-certainty as feedback. These approaches report large jumps on Qwen2.5-Math-7B, occasionally matching or surpassing stronger supervised baseline. Other variants explore noisy (Lv et al., 2025) or even random rewards (Shao et al., 2025) for Qwen2.5. However, these methods fails to transfer to Llama or OLMo (OLMo et al., 2025), suggesting model-specific idiosyncrasies.
@@ 2.2 Factors Influencing Performance on Math
The choice of pretraining corpora plays a crucial role in shaping the reasoning abilities of LLMs, particularly in mathematical domains. Several math-specific datasets (Paster et al., 2024; Han et al., 2024; Wang et al., 2024; Allal et al., 2025; Stojanovski et al., 2025) have been proposed and shown to significantly enhance performance on relevant benchmarks. Further, Wang et al. (2025b) finds that mid-training Llama models on high-quality mathematical corpora substantially improve their capacity, both at the base level and after reinforcement learning. However, evaluation of math capability can be misleading due to potential test data contamination. Xu et al. (2024) observed that certain widely-used benchmarks may be partially included in the pretraining corpus of LLMs, including early versions of Qwen. On the other hand, Liu et al. (2025) found that omitting dialogue-style prompting can lead to improved mathematical reasoning. These observations underscore the presence of multiple confounding factors in evaluating model performance, motivating the need for rigorous and systematic analysis.
[…]
@@ 3.1 Model Selection
Prior researches on mathematical reasoning with LLMs focus predominantly on the Qwen-2.5 (Yang et al., 2024a; b). Accordingly, we center our study on four representative checkpoints from this series: Qwen2.5-7B, Qwen2.5-7B-Instruct, Qwen2.5-Math-7B, and Qwen2.5-Math-7B-Instruct. For a controlled comparison, we also evaluate Llama3.1-8B and Llama3.1-8B-Instruct (Dubey et al., 2024), which possess comparable parameter counts and thus help isolate model-specific differences in behavior.
@@ 3.2 Evaluation of Memorization Capability
Specifically, we prompt the model to complete the remaining parts of a problem based on partial prefixes. To evaluate its performance, we use the Partial-Prompt Completion Rate measured by ROUGE-L (Lin, 2004), which calculates the overlap of the longest common subsequence between the generated and reference text, capturing fluency and sentence-level structure. Additionally, we utilize Exact Match (EM) accuracy, which checks if the model’s output exactly matches the reference. The final EM score is the average across all instances. Higher EM indicates a greater proportion of partial-prompts that model can recall exactly.
@@ 3.3 RLVR-Based Evaluation
Group Relative Policy Optimization (GRPO) (Shao et al., 2024) is adopted as our RLVR algorithm. Formally, for each question $q$, GRPO samples a group of outputs $\{o_{1},\cdots,o_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$ and then optimizes the policy model by maximizing the following objective:
$\begin{aligned} \mathcal{J}&{}_{\text{GRPO}}(\theta)=\mathbb{E}_{q,\{o_{i}\}_{i=1}^{G}}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left\{\min\left(r_{i,t}\hat{A}_{i,t},\text{clip}\left(r_{i,t},1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\right)-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}||\pi_{\text{ref}}\right]\right\},\end{aligned}$
@@
Following Shao et al. (2025), we consider following spurious reward types for RLVR:
@@ 4.1 Spurious Rewards on MATH-500
Following the work of Shao et al. (2025), we replicate the performance of Qwen2.5-Math-7B and Llama3.1-8B-Instruct on the MATH-500 benchmark under various reward signal configurations, using the same experimental setup. The accuracy curves of MATH-500 are shown in Fig. 3. Interestingly, the results demonstrate that while random rewards and mv-incorrect rewards noticeably boost accuracy for Qwen2.5-Math-7B, they have little or even adverse impacts on the performance of Llama3.1-8B-Instruct. Additionally, we apply the same RLVR procedure to Qwen2.5-Math-7B-Instruct and discover that the resulting gains are marginal when compared with those of Qwen2.5-Math-7B, indicating that the two Qwen variants exhibit differential sensitivity to RLVR under spurious rewards.
Considering the base and instruct variants of Qwen are trained under different paradigms: the former is pre-trained as a general language model without exposure to any dialogue-specific templates, whereas the latter undergoes an additional instruction-tuning stage on data wrapped in a fixed dialogue template. This mismatch creates a training–testing gap for the base model at the start point of RLVR, and its initial accuracy is therefore likely underestimated. Consequently, to obtain a fair estimate of each model’s starting point, we next measure performance under four decoding configurations as shown in Tab. 1. The corresponding results are summarized in Fig. 4. Surprisingly, we discover that applying the official chat template substantially degrades performance for the Qwen base model: both Qwen-2.5-7B and Qwen-2.5-Math-7B suffer pronounced drops once the template is enabled.
Table 1: The sampling parameters used under different generation configurations. Greedy (w/o Template) refers to greedy decoding performed with the default model.generate(…) call. Avg@16 (w/o Template) indicates that we sample 16 candidate answers with vLLM and report their average. Greedy (w/ Template) denotes the use of the official chat template combined with the default greedy decoding, whereas Avg@16 (w/ Template) applies the official chat template together with the average over 16 rollout samples. In all configurations, the max-tokens parameter is set to 4096.
Configuration Do-Sample Temperature Top-P Top-K Chat Template
Greedy (w/o Template) False 1.0 1.0
Avg@16 (w/o template) 0.7 0.8 20
Greedy (w/ Template) False 1.0 1.0
Avg@16 (w/ template) 0.7 0.8 20
@@
Building on this observation, we report two additional initial accuracy for reference in Fig. 3: (i) Greedy (w/o Template) corresponds to the best performance of the initial model, and (ii) pass@16, adopted from Yue et al. (2025), serves as a plausible performance upper bound for the initial model. Viewed against these baselines, the seeming ‘RL gains’ of Qwen-Math-7B largely reflect adaptation to the template format and merely converge to the Greedy (w/o Template) baseline, indicative of memory recall rather than genuine mathematical generalization. However, spurious rewards, e.g., random and mv-incorrect, still boost the accuracy of Qwen base and maintain the performance of Qwen instruct, while degrading Llama eventually. This is the point we need to further explore in the following sections.
Table 2: Accuracy (Exact Match, EM) and ROUGE-L scores on several datasets (lower scores in gray) under different prompt prefix ratios in greedy decoding mode without applying chat template, namely Greedy (w/o Template) configuration.
[…]

### v3

@@ Abstract
Reasoning in large language models has long been a central research focus, and recent studies employing reinforcement learning (RL) have introduced diverse methods that yield substantial performance gains with minimal or even no external supervision. Surprisingly, some studies even suggest that random or incorrect reward signals can enhance performance. However, these breakthroughs are predominantly observed for the mathematically strong Qwen2.5 series on benchmarks such as MATH-500, AMC, and AIME, and seldom transfer to models like Llama, which warrants a more in-depth investigation. In this work, our empirical analysis reveals that pre-training on massive web-scale corpora leaves Qwen2.5 susceptible to data contamination in widely used benchmarks. Consequently, conclusions derived from contaminated benchmarks on Qwen2.5 series may be unreliable. To obtain trustworthy evaluation results, we introduce a generator that creates fully clean arithmetic problems of arbitrary length and difficulty, dubbed RandomCalculation. Using this leakage-free dataset, we show that only accurate reward signals yield steady improvements that surpass the base model’s performance boundary in mathematical reasoning, whereas random or incorrect rewards do not. Moreover, we conduct more fine-grained analyses to elucidate the factors underlying the different performance observed on the MATH-500 and RandomCalculation benchmarks. Consequently, we recommend that future studies evaluate models on uncontaminated benchmarks and, when feasible, test various model series to ensure trustworthy conclusions about RL and related methods.
Code: https://github.com/wumingqi/LLM-Math-Evaluation
@@ 1 Introduction
In recent years, advances in reinforcement learning (RL) have markedly strengthened the reasoning abilities of large language models (LLMs). Flagship systems, including OpenAI’s o1 (openai-o1; openai-o3; openai-gpt4o), DeepSeek-R1 (DeepSeek-R1), and QwQ (qwq; qwq32b), already match or exceed human-level accuracy on a variety of challenging benchmarks. Among open-source contenders that excel in these mathematical benchmarks, the Qwen family (DBLP:qwen2.5-tech-report; DBLP:qwen2.5-math-tech-report; DBLP:Qwen3-Technical-Report), ranging from 0.5B to 72B and pre-trained on up to 36T high-quality tokens, produces state-of-the-art results in language understanding, mathematics, programming, and preference alignment.
Within this landscape, mathematical reasoning emerges as a particularly discriminative test bed because it demands precise symbolic manipulation and multi-step logical deduction. Standard suites such as MATH-500 (DBLP:conf/nips/HendrycksBKABTS21), AIME (numina_math), AMC (numina_math), and Minerva Math (DBLP:minerva_math) require models to parse natural-language problem statements, uncover the latent mathematical structure, and generate exact numeric answers. Recent work has further enhanced this capability by reinforcement learning with verifiable rewards (RLVR) (DeepSeek-R1): rule-based reward that returns 1 when the predicted answer equals ground truth and 0 otherwise. Because the reward is computed analytically, RLVR removes the need for a separate learned reward model, lowering computational cost while providing a precise training signal, especially attractive for domains like mathematics, where solutions are unambiguous.
@@
Qwen2.5-7B Output it true that $2^{24}$ is a perfect $n^{\text{th}}$ power? To determine for how many positive integers $n>1$ the number $2^{24}$ is a perfect $n$-th power… Let’s count these values. There are 77 such values… The final answer is $\boxed{7}$.
Llama3.1-8B Output the set of positive divisors of $n$ (including 11 and $n$) less than 5050 ? (A) 15 (B) 16 (C) 17 … The number of positive divisors of $n$ is given by… The number of factors of each of these numbers is… The answer is $\boxed{\textbf{(B) }16}$
@@
Although RL nominally depends on accurate reward signals to guide training, recent studies (TTRL; Spurious-Rewards) find that even random or incorrect rewards can improve Qwen’s performance on standard math benchmarks, while the same procedures offer little or no benefit to Llama3.1-8B (llama3). To understand why these seemingly problem-agnostic rewards help Qwen but not Llama, we undertake a systematic comparison of two model families under identical training protocols. We consider two working hypotheses to explain this phenomenon. (i) Data Contamination: Considering Qwen2.5 is pre-trained on massive web-scale corpora from the Internet, including GitHub repositories that store benchmark problems alongside their official solutions. If segments of the evaluation benchmarks leaked into the pre-training corpus, spurious rewards could cue the model to retrieve memorized answers rather than acquire new reasoning skills. (ii) Strong Math Capacity: Qwen’s pre-training endows it with better mathematical capacity than Llama, so even noisy policy-gradient updates appear to help on MATH-500. However, if strong capacity is the real driver, the same spurious rewards should still work on a clean benchmark. Distinguishing between these possibilities requires both a leakage audit and a rigorously out-of-distribution RLVR evaluation.
To assess the extent of potential data contamination in popular mathematical benchmarks, we propose two metrics: partial-prompt completion rate (can the model reconstruct the tail of a problem?) and partial-prompt answer accuracy (can the model give the correct answer with an incomplete problem?). As shown in Fig. 1, Qwen can indeed complete the problem accurately and provide the correct answer, whereas Llama fails. Furthermore, prompting with the first 60% of each MATH-500 problem, Qwen2.5-Math-7B regenerates the remaining 40% with a 54.60% exact-match rate and answers 53.6% of these incomplete problems correctly. In contrast, Llama3.1-8B scores 3.8% and 2.4% on both metrics. Crucially, on the newly released LiveMathBench (version 202505) (DBLP:LiveMathBench), Qwen’s completion rate drops sharply to 0.0%, consistent with Llama’s 0.0%. Its partial-prompt answer accuracy also falls to just 2.0%, comparable to Llama’s 1.0%. These results confirm that Qwen’s pre-training corpus suffers from test data contamination. So, results derived from MATH-500 and similar datasets for Qwen should be interpreted with caution.
@@ 2 Related Works
A growing body of work investigates how reinforcement learning (RL) can amplify the mathematical-reasoning capacity of the open-source Qwen2.5 family. Early studies use verifiable rewards that score an answer as 1 / 0 by exact numerical agreement. Test-time RL (TTRL) applies this signal on-the-fly during inference and yields sizeable gains on MATH-500 and AIME2024. Subsequent efforts pursue extreme data efficiency: few labeled (1-shot-RL; Few-Shot-RL) or even unlabeled examples (one-shot-EM; Absolute-Zero) can suffice to boost performance. A parallel research replaces external supervision with intrinsic signals derived from the model itself: RENT; EM-RL reward low-entropy output distributions, while INTUITOR; Self-Train rely on self-consistency or self-certainty as feedback. These approaches report large jumps on Qwen2.5-Math-7B, occasionally matching or surpassing stronger supervised baseline. Other variants explore noisy (lv-PRP) or even random rewards (Spurious-Rewards) for Qwen2.5. However, these methods fails to transfer to Llama or OLMo (DBLP:olmo), suggesting model-specific idiosyncrasies.
@@ 2.2 Factors Influencing Performance on Math
The choice of pretraining corpora plays a crucial role in shaping the reasoning abilities of LLMs, particularly in mathematical domains. Several math-specific datasets (OpenWebMath; InfiMM-WebMath; MathPile; SmolLM2) have been proposed and shown to significantly enhance performance on relevant benchmarks. Further, OctoThinker finds that mid-training Llama models on high-quality mathematical corpora substantially improve their capacity, both at the base level and after reinforcement learning. However, evaluation of math capability can be misleading due to potential test data contamination. DBLP:Benchmarking-Benchmark observed that certain widely-used benchmarks may be partially included in the pretraining corpus of LLMs, including early versions of Qwen. On the other hand, DBLP:Understanding-R1 found that omitting dialogue-style prompting can lead to improved mathematical reasoning. These observations underscore the presence of multiple confounding factors in evaluating model performance, motivating the need for rigorous and systematic analysis.
@@ 3.1 Model Selection
Prior researches on mathematical reasoning with LLMs focus predominantly on the Qwen-2.5 (DBLP:qwen2.5-tech-report; DBLP:qwen2.5-math-tech-report). Accordingly, we center our study on four representative checkpoints from this series: Qwen2.5-7B, Qwen2.5-7B-Instruct, Qwen2.5-Math-7B, and Qwen2.5-Math-7B-Instruct. For a controlled comparison, we also evaluate Llama3.1-8B and Llama3.1-8B-Instruct (llama3), which possess comparable parameter counts and thus help isolate model-specific differences in behavior.
@@ 3.2 Evaluation of Memorization Capability
Specifically, we prompt the model to complete the remaining parts of a problem based on partial prefixes. To evaluate its performance, we use the Partial-Prompt Completion Rate measured by ROUGE-L (lin-2004-rouge), which calculates the overlap of the longest common subsequence between the generated and reference text, capturing fluency and sentence-level structure. Additionally, we utilize Exact Match (EM) accuracy, which checks if the model’s output exactly matches the reference. The final EM score is the average across all instances. Higher EM indicates a greater proportion of partial-prompts that model can recall exactly.
@@ 3.3 RLVR-Based Evaluation
Group Relative Policy Optimization (GRPO) (DeepSeekMath) is adopted as our RLVR algorithm. Formally, for each question $q$, GRPO samples a group of outputs $\{o_{1},\cdots,o_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$ and then optimizes the policy model by maximizing the following objective:
[…]
$\left.\left.\text{clip}\left(r_{i,t},1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\right)-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}||\pi_{\text{ref}}\right]\right\},$
@@
Following Spurious-Rewards, we consider following spurious reward types for RLVR:
@@ 4.1 Spurious Rewards on MATH-500
Following the work of Spurious-Rewards, we replicate the performance of Qwen2.5-Math-7B and Llama3.1-8B-Instruct on the MATH-500 benchmark under various reward signal configurations, using the same experimental setup. The accuracy curves of MATH-500 are shown in Fig. 3. Interestingly, the results demonstrate that while random rewards and mv-incorrect rewards noticeably boost accuracy for Qwen2.5-Math-7B, they have little or even adverse impacts on the performance of Llama3.1-8B-Instruct. Additionally, we apply the same RLVR procedure to Qwen2.5-Math-7B-Instruct and discover that the resulting gains are marginal when compared with those of Qwen2.5-Math-7B, indicating that the two Qwen variants exhibit differential sensitivity to RLVR under spurious rewards.
Considering the base and instruct variants of Qwen are trained under different paradigms: the former is pre-trained as a general language model without exposure to any dialogue-specific templates, whereas the latter undergoes an additional instruction-tuning stage on data wrapped in a fixed dialogue template. This mismatch creates a training–testing gap for the base model at the start point of RLVR, and its initial accuracy is therefore likely underestimated. Consequently, to obtain a fair estimate of each model’s starting point, we next measure performance under four decoding configurations as shown in Tab. 1. The corresponding results are summarized in Fig. 4. Surprisingly, we discover that applying the official chat template substantially degrades performance for the Qwen base model: both Qwen2.5-7B and Qwen2.5-Math-7B suffer pronounced drops once the template is enabled.
Configuration # Sample Temp. Top-P Top-K
Greedy (w/o Template) 1 1.0 1.0 1
Avg@16 (w/o template) 16 0.7 0.8 20
Greedy (w/ Template) 1 1.0 1.0 1
Avg@16 (w/ template) 16 0.7 0.8 20
Table 1: The sampling parameters used under different generation configurations. Greedy sampling is performed using the default model.generate(…) function, while random sampling is implemented using vLLM. w/ Template indicates the usage of official chat template.
@@
Building on this observation, we report two additional initial accuracy for reference in Fig. 3: (i) Greedy (w/o Template) corresponds to the best performance of the initial model, and (ii) pass@16, adopted from DBLP:DoesRLIncentivizeReasoning, serves as a plausible performance upper bound for the initial model. Viewed against these baselines, the seeming ‘RL gains’ of Qwen2.5-Math-7B largely reflect adaptation to the template format and merely converge to the Greedy (w/o Template) baseline, indicative of memory recall rather than genuine mathematical generalization. However, spurious rewards, e.g., random and mv-incorrect, still boost the accuracy of Qwen base and maintain the performance of Qwen instruct, while degrading Llama eventually. This is the point we need to further explore in the following sections.
Model Dataset Size 80%-Problem 60%-Problem 40%-Problem
ROUGE-L EM ROUGE-L EM ROUGE-L EM
Qwen2.5-Math-7B MATH-500 500 81.25 65.80 78.06 54.60 69.01 39.20
AMC 83 77.38 55.42 70.25 42.17 75.17 36.14
AIME2024 30 74.04 56.67 55.31 20.00 57.72 16.67
AIME2025 30 54.71 16.67 34.88 0.00 27.43 0.00
MinervaMath 272 36.08 2.94 31.22 0.37 29.35 0.00
LiveMathBench 100 42.76 5.00 32.78 0.00 29.97 0.00
Qwen2.5-7B MATH-500 500 66.42 40.20 60.98 21.20 50.36 8.20
AMC 83 73.24 49.40 64.42 33.73 63.79 28.92
AIME2024 30 59.80 30.00 48.69 13.33 44.65 10.00
AIME2025 30 54.61 10.00 37.59 0.00 30.30 0.00
MinervaMath 272 35.24 2.94 32.35 0.37 27.89 0.00
LiveMathBench 100 41.15 4.00 32.74 0.00 27.95 0.00
Llama3.1-8B MATH-500 500 48.33 17.80 40.55 3.80 32.07 0.60
AMC 83 44.54 4.82 30.62 0.00 27.10 0.00
AIME2024 30 50.50 13.33 30.80 0.00 26.08 0.00
AIME2025 30 47.04 10.00 33.49 0.00 25.20 0.00
MinervaMath 272 36.24 2.21 29.52 0.00 27.11 0.00
LiveMathBench 100 35.55 5.00 31.93 0.00 26.88 0.00
Table 2: Accuracy (EM) and ROUGE-L on several datasets (lower scores in gray) under different prompt prefix ratios with Greedy (w/o Template) configuration. Suspicious accuracy is highlighted in bold.